# Concentration and fluctuations of sine–Gordon measure around topological multi-soliton manifold

### v1

@@ 6.1. Free energy
$\Theta^{0}:=\int_{0}^{1}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta^{0}(t)dt=-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}.$ (6.11)
In (6.10) the choice of the drift $\theta^{0}(t)$ is admissible $\theta^{0}\in\mathbb{H}_{a}$ since it is an adapted process due to the cutoff $\mathbf{1}_{\{t>1-\eta\}}$ and it satisfies the required regularity condition thanks to the truncation to the modes $1\leq n\leq N$. By plugging (6.10) and (6.11) into (6.9), we obtain
$\varepsilon\log Z_{\varepsilon}^{Q}\geq\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))dx-\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\bigg].$ (6.12)
@@
$\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt=\frac{1}{2}\|(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\eta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\|_{L^{2}_{x}}^{2}.$
@@
$\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\leq\frac{1+\delta}{2}\|\partial_{x}m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon}\|_{L^{2}}^{2}+\frac{C_{\delta}}{2}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\eta))\|_{L^{2}}^{2}+\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}\big)$ (6.13)
@@
$\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))$
$=\cos(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})-\sin(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})\cdot\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))+O\big((\sqrt{\varepsilon}(Y-Y_{N}(1-\eta)))^{2}\big).$ (6.14)
@@
$\hphantom{X}-\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))|^{2}dx\bigg]$
$\hphantom{X}-C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}-C_{\delta}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\eta))\|_{L^{2}}^{2},$ (6.15)
where we used $\mathbb{E}\big[\sqrt{\varepsilon}(Y-Y_{N}(1-\eta))\big]=0$. Using Lemma 2.2 and $\partial_{x}\ell^{Q}=\frac{\pi Q}{L_{\varepsilon}}$, we obtain
@@
$\hphantom{X}-c\varepsilon L_{\varepsilon}^{2}-C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}-C_{\delta}\varepsilon\|\partial_{x}Y_{N}(1-\eta)\|_{L^{2}}^{2}.$
@@
This completes the proof of Proposition 6.1.
In order to prove the main result of this section, that is, the large deviation estimates, we need the following lemma.
@@
Let $Q\in\mathbb{Z}$ with $Q>0$ and $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$. Then the following holds.
  • (i)
    Let $\delta>0$. Then we have
    $\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi).$
  • (ii)
    Let $d,\delta>0$. Then
    $\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\end{subarray}}E(\phi).$
@@
As in (6.1), using the Boué–Dupuis formula (Lemma 2.1) and the change of variables $\sqrt{\varepsilon}\theta\to\theta$, we write
$\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta\}}\bigg]$
$\leq\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\cdot\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta\}}\Big\}\bigg]$
$\leq\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx\cdot\mathbf{1}_{\{\textup{dist}(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta,\mathcal{M}^{Q})\geq\delta\}}-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg].$ (6.16)
By taking the change of variable $W=\ell^{Q}+\Theta$, we write
[…]
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg]$
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[\bigg(-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+W))dx-\frac{1}{2}\|\partial_{x}W\|_{L^{2}}^{2}\bigg)\cdot\mathbf{1}_{\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q})\geq\delta\}}$
$\hphantom{XXXXXXXXXXXX}-\frac{1}{2}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}+\int_{-L_{\varepsilon}}^{L_{\varepsilon}}\partial_{x}W\cdot\partial_{x}\ell^{Q}dx\bigg].$ (6.17)
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, we have $\mathbb{E}\big[\|\sqrt{\varepsilon}Y\|_{L^{2}}^{2}\big]\to 0$ by (2.8). Therefore, with high probability
$\{\textup{dist}(W,\mathcal{M}_{Q})\geq 2\delta\}\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q})\geq\delta\}.$ (6.18)
@@
$\hphantom{XXXXXXXXX}+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\eta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}$ (6.19)
for any $\eta>0$ and $\kappa>0$. This implies that under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\eta\to 0$, and $\kappa\to 0$ in order, we obtain
$\limsup_{\varepsilon\to 0}\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi).$
This proves the first part of Lemma 6.3.
For the second part of Lemma 6.3, we only need to modify the following part. Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, we have $\mathbb{E}\big[\|\sqrt{\varepsilon}Y\|_{L^{2}}^{2}\big]\to 0$ by (2.8). Therefore, with high probability
$\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\}$ $\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q}^{<d})<\delta\}.$ (6.20)
By proceeding as in obtaining (6.19), together with (6.20), we obtain
$\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{<d}_{Q})<\delta\}}\bigg]$
$\leq\sup_{W\in\ell^{Q}+\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[\bigg(-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\kappa)(1-\cos W)dx-\frac{1-\eta}{2}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx\bigg)\mathbf{1}_{\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\}}\bigg]$
$\hphantom{XXXXXXXXX}+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\eta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}.$
Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\eta\to 0$, and $\kappa\to 0$ in order, we obtain the desired result.
6.2. Large deviations
In this subsection, we study the concentration of the measure $\rho_{\varepsilon}^{Q}$.
@@ 6.2. Large deviations
Note that
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)=\frac{\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]}{Z_{\varepsilon}^{Q}}.$
This implies
$\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)$
$=\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\Big\}\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\bigg]-\varepsilon\log Z_{\varepsilon}^{Q}.$ (6.21)
@@
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$ (6.22)
@@
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2}.$ (6.23)
This completes the proof of (1.6).
We now prove (1.8). By following the argument used to obtain (6.22) with Lema 6.3, we get
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-\inf_{\begin{subarray}{c}\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$ (6.24)
Under the condition $\frac{1}{2}e^{-d}\geq c\cdot\delta$, we can apply Lemma 3.10 to obtain
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d}.$ (6.25)
This completes the proof of Theorem 1.2.
Remark 6.4.
@@
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta\}\big)\leq-c\delta^{2},$ (6.26)
where $\mathcal{M}_{Q}^{\varepsilon}$ is defined as in (4.8), and under $\frac{1}{2}e^{-d}\geq c\cdot\delta$
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d})<\delta\}\big)\leq-ce^{-d},$ (6.27)
where $\mathcal{M}^{\varepsilon,<d}_{Q}$ is defined as in (4.9). Therefore, from (6.26), (6.27) and $\frac{1}{2}e^{-d}\geq c\cdot\delta$, by choosing $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\frac{1}{\varepsilon}}\big|$, we have
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta_{\varepsilon}\}\big)\lesssim e^{-c\log\frac{1}{\varepsilon}}$ (6.28)
$\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}$ (6.29)
for $0<\varepsilon<\varepsilon_{0}$, where $\varepsilon_{0}$ is a small number.
Remark 6.5.
Note that
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d})<\delta\}\subset\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}$ (6.30)
and
$\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d})<\delta\}$
$\subset\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d})<\delta\big\},$ (6.31)
where
$\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d}$
$:=\Big\{\sum_{j=1}^{Q}m^{\varepsilon}(\cdot-\xi_{j}):-\overline{L}_{\varepsilon}\leq\xi_{1}\leq\dots\leq\xi_{Q}\leq\overline{L}_{\varepsilon}\;\;\textup{and}\;\;d\leq\min_{i\neq j}|\xi_{i}-\xi_{j}|<(1+\eta)d\Big\}.$
By following the proof of Theorem 1.2 together with Lemma 3.13 and $\frac{1}{2}e^{-d}\geq c\cdot\delta$,
$\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d})<\delta\big\}\Big)$
$\leq\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,d\leq\cdot<(1+\eta)d})<\delta\big\}\Big)\leq-c\cdot e^{-(1+\eta)d}.$
Since the distance $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and the collision scale $d_{\varepsilon}=\big|\log\sqrt{\varepsilon\log\tfrac{1}{\varepsilon}}\big|$, chosen above in (6.29), satisfy $\frac{1}{2}e^{-d_{\varepsilon}}\geq c\cdot\delta_{\varepsilon}$, we have
$\rho_{\varepsilon}^{Q}\Big(\big\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq d})<\delta\}\setminus\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\geq(1+\eta)d})<\delta\big\}\Big)\lesssim\exp\Big\{-\frac{c}{\varepsilon}\cdot e^{-(1+\eta)d_{\varepsilon}}\Big\}\lesssim e^{-c\cdot\varepsilon^{-\frac{1}{2}+}}.$ (6.32)
Therefore, most of the probability mass is concentrated on the set $\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,\,\geq(1+\eta)d_{\varepsilon}})<\delta_{\varepsilon}\}$. For the use of this large deviation estimate, see Remark 7.2 and Lemma 7.1.
7. Preliminaries for the central limit theorem
7.1. Partition into a typical event and large deviation events
Using the large-deviation estimates established above, we decompose the entire space into a typical event and large-deviation events as follows.

### v2

@@ 5. Gaussian measures associated with Schrödinger operators
(6.16) $\Theta^{0}:=\int_{0}^{1}(-\partial_{x}^{2})^{-\frac{1}{2}}\theta^{0}(t)dt=-\sqrt{\varepsilon}Y_{N}(1-\zeta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}.$
In (6.15) the choice of the drift $\theta^{0}(t)$ is admissible $\theta^{0}\in\mathbb{H}_{a}$ since it is an adapted process due to the cutoff $\mathbf{1}_{\{t>1-\zeta\}}$ and it satisfies the required regularity condition thanks to the truncation to the modes $1\leq n\leq N$. By plugging (6.15) and (6.16) into (6.14), we obtain
(6.17) $\varepsilon\log Z_{\varepsilon}^{Q}\geq\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))dx-\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\bigg].$
@@
$\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt=\frac{1}{2}\|(-\partial_{x}^{2})^{\frac{1}{2}}(-\sqrt{\varepsilon}Y_{N}(1-\zeta)-\ell^{Q}+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}})\|_{L^{2}_{x}}^{2}.$
@@
(6.18) $\frac{1}{2}\int_{0}^{1}\|\theta^{0}(t)\|_{L^{2}_{x}}^{2}dt\leq\frac{1+\delta}{2}\|\partial_{x}m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon}\|_{L^{2}}^{2}+\frac{C_{\delta}}{2}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\zeta))\|_{L^{2}}^{2}+\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}\big)$
@@
$\cos(\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta))+m^{\varepsilon}_{\xi_{1},\dots,\xi_{Q}}))$
(6.19) $=\cos(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})-\sin(m_{\xi_{1},\dots,\xi_{Q}}^{\varepsilon})\cdot\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta))+O\big((\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta)))^{2}\big).$
@@
$\hphantom{X}-\mathbb{E}\bigg[\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta))|^{2}dx\bigg]$
(6.20) $\hphantom{X}-C_{\delta}\|\partial_{x}\ell^{Q}\|_{L^{2}}^{2}-C_{\delta}\big(\|\partial_{x}(\sqrt{\varepsilon}Y_{N}(1-\zeta))\|_{L^{2}}^{2},$
where we used $\mathbb{E}\big[\sqrt{\varepsilon}(Y-Y_{N}(1-\zeta))\big]=0$. Using Lemma 2.2 and $\partial_{x}\ell^{Q}=\frac{\pi Q}{L_{\varepsilon}}$, we obtain
@@
$\hphantom{X}-c\varepsilon L_{\varepsilon}^{2}-C_{\delta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}-C_{\delta}\varepsilon\|\partial_{x}Y_{N}(1-\zeta)\|_{L^{2}}^{2}.$
@@
Let $Q\in\mathbb{Z}$ with $Q>0$ and $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$. Then for $d,\delta>0$,
$\limsup_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\Big]$ $\leq\;\;-\!\!\!\!\!\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}\!\!\!E(\phi),$
$\limsup_{\varepsilon\to 0}\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta}\Big]$ $\leq\;\;-\!\!\!\!\!\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\end{subarray}}\!\!\!E(\phi).$
@@
Note that
$\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}}\Big]\leq\varepsilon\log\mathbb{E}_{\mu^{Q}_{\varepsilon}}\bigg[\exp\Big\{-\frac{1}{\varepsilon}\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos\phi)dx\cdot\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}^{Q})\geq\delta\}}\Big\}\bigg]$
since the indicator may only take values in $\{0,1\}$. Proceeding as in the previous proof, we can bound the above expression by
(6.21) $\sup_{\Theta\in\dot{\mathbb{H}}^{1}}\mathbb{E}\bigg[-\int_{-L_{\varepsilon}}^{L_{\varepsilon}}(1-\cos(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta))dx\cdot\mathbf{1}_{\{\textup{dist}(\sqrt{\varepsilon}Y+\ell^{Q}+\Theta,\mathcal{M}^{Q})\geq\delta\}}-\frac{1}{2}\|\partial_{x}\Theta\|_{L^{2}}^{2}\bigg].$
Take the change of variable $W=\ell^{Q}+\Theta$. Under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, we have $\mathbb{E}\big[\|\sqrt{\varepsilon}Y\|_{L^{2}}^{2}\big]\to 0$ by (2.8). Therefore, with high probability
(6.22) $\{\textup{dist}(W,\mathcal{M}_{Q})\geq 2\delta\}\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q})\geq\delta\}.$
@@
(6.23) $\hphantom{XXXXXXXXX}+C_{\kappa}\varepsilon L_{\varepsilon}^{2}+C_{\eta}\frac{(\pi Q)^{2}}{L_{\varepsilon}}$
for any $\eta>0$ and $\kappa>0$. Here we have multiplied $\int_{-L_{\varepsilon}}^{L_{\varepsilon}}|\partial_{x}W|^{2}dx$ by the indicator as it is positive. This implies that under the condition $L_{\varepsilon}=\varepsilon^{-\frac{1}{2}+}$, by taking the limits $\varepsilon\to 0$, $\eta\to 0$, and $\kappa\to 0$ in order, we obtain the first part of Lemma 6.2.
For the second part of Lemma 6.2, the only modification is that now with high probability
(6.24) $\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\}$ $\subset\{\textup{dist}(\sqrt{\varepsilon}Y+W,\mathcal{M}_{Q}^{<d})<\delta\}.$
So for $\varepsilon\log Z_{\varepsilon}^{Q}\Big[\mathbf{1}_{\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta}\Big]$ we get the same bound as in (6.23), with the indicator replaced by $\mathbf{1}_{\{\textup{dist}(W,\mathcal{M}_{Q}^{<d})<\frac{\delta}{2}\}}$. So we again obtain the desired result.
@@
(6.25) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-\inf_{\begin{subarray}{c}\phi\in\mathcal{C}_{Q}\\ \textup{dist}(\phi,\mathcal{M}_{Q})\geq 2\delta\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
@@
(6.26) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q})\geq\delta\}\big)\leq-c\delta^{2}.$
This proves (1.6). Moreover, by (6.2), Proposition 6.1 and Lemma 6.2
(6.27) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-\inf_{\begin{subarray}{c}\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\tfrac{\delta}{2}\end{subarray}}E(\phi)+\inf_{\phi\in\mathcal{C}_{Q}}E(\phi).$
Under the condition $ce^{-d}\geq\delta^{2}$, we can apply Lemma 3.10 to obtain
(6.28) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{<d})<\delta\big)\leq-ce^{-d}.$
This proves (1.8), thus completing the proof of Theorem 1.2.
Corollary 6.3.
Under $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, we have
(6.29) $\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta_{\varepsilon}\}\big)\lesssim e^{-c\log\frac{1}{\varepsilon}}$
(6.30) $\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d_{\varepsilon}})<\delta_{\varepsilon}\}\big)\lesssim e^{-c\log\frac{1}{\varepsilon}}$
where $\mathcal{M}_{Q}^{\varepsilon}$, $\mathcal{M}^{\varepsilon,<d}_{Q}$ are approximating soliton manifolds, defined as in (4.8) and (4.9).
Proof.
@@
(6.31) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon})\geq\delta\}\big)$ $\leq-c\delta^{2}$
(6.32) $\limsup_{\varepsilon\to 0}\varepsilon\log\rho_{\varepsilon}^{Q}\big(\{\textup{dist}(\phi,\mathcal{M}_{Q}^{\varepsilon,<d})<\delta\}\big)$ $\leq-ce^{-d},$
where the condition $c\cdot e^{-d}\geq\delta^{2}$ follows from Lemma 3.10. Therefore, from (6.31), (6.32) and $c\cdot e^{-d}\geq\delta^{2}$, by choosing $\delta_{\varepsilon}=\eta\sqrt{\varepsilon\log\frac{1}{\varepsilon}}$ and $d_{\varepsilon}=\big|\log(\varepsilon\log\frac{1}{\varepsilon})\big|$, we obtain (6.29) and (6.30).
7. Change of coordinates
The main result of this section is Proposition 7.2 which allows us to turn the integration over $\phi$ to an integration over $(\xi_{1},\cdots,\xi_{Q})$ and $v\in V_{\xi_{1},\dots,\xi_{Q}}$, up to small errors. To this end we first use the large-deviation estimates established above to decompose the entire space into a typical event and large-deviation events as follows. Recall the definition of $\mathcal{M}^{\varepsilon,\geq d}_{Q}$ in (4.10).